A petroleum company has two different sources of crude oil. The first source provides crude oil that is 20% hydrocarbons, and the second one provides crude oil that is 60% hydrocarbons. In order to obtain 160 gallons of crude oil that is 45% hydrocarbons, how many gallons of crude oil must be used from each of the two sources?
Answer by TimothyLamb(4379)

x = gallons of 20%
y = gallons of 60%
---
20x + 60y = 45*160
x + y = 160
---
put the system of linear equations into standard form
---
20x + 60y = 7200
x + y = 160
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = gallons of 20% = 60
y = gallons of 60% = 100
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php